OpenStudy (anonymous):

is 4 × 4 × 4 × 4 × 4=4^4

OpenStudy (anonymous):

idk if its 4 ^ 4 or 4 ^ 5 :O

OpenStudy (anonymous):

4^5

OpenStudy (anonymous):

So you also count the first 4 ? Because there are 5, 4's there and i thought you might count the first 4 as the 4...in 4^4

OpenStudy (anonymous):

Yep you count all the 4s, the number of times it appears is the exponent, the 4 is the base

OpenStudy (anonymous):

like 4 = 1 to the 4th poewr so the other 4's would =4 so 4^4???

OpenStudy (anonymous):

yeah like five 4's right? so one 4 is the base then the other 4's are the 4 Right?!\[4^{4}\]

OpenStudy (anonymous):

The base will be 4, the number of times 4 appears is the exponent. So 4^5

OpenStudy (anonymous):

You don't take away one of the 4s because you use it as the base

OpenStudy (anonymous):

D: How? I'm not understanding!! *sniff sniff* i'm so dumb...

OpenStudy (anonymous):

Don't worry, you'll get it

OpenStudy (anonymous):

But how if I don't understand it :(((

OpenStudy (anonymous):

4^1=4 4^2=4*4 4^3=4*4*4 4^4=4*4*4*4 4^5=4*4*4*4*4

OpenStudy (anonymous):

:D....Heheheeeeeee

OpenStudy (anonymous):

My light bulb just clicked :P

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

So example 10 × 10 × 10 × 10 × 10 Would be 10^5? :D

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

YAY!

OpenStudy (anonymous):

And 2 × 2 × 2 × 2 × 2 × 2 is 2^6 :D?

OpenStudy (anonymous):

that's right